# Chain-linked Multiple Matrix Integration via Embedding Alignment

### v2

@@ 1 Introduction
The development of large-scale data collection and sharing has sparked considerable research interests in integrating data from diverse sources to efficiently uncover underlying signals. This problem is especially pertinent in fields such as healthcare research (Hong et al., 2021; Zhou et al., 2023), genomic data integration (Maneck et al., 2011; Tseng et al., 2015; Cai et al., 2016), single-cell data integration (Stuart et al., 2019; Ma et al., 2024), and chemometrics (Mishra et al., 2021). In this paper we consider a formulation of the problem where each source $i$ corresponds to a partially observed submatrix $\mathbf{M}^{(i)}$ of some matrix $\mathbf{M}$, and the goal is to integrate these $\{\mathbf{M}^{(i)}\}$ to recover $\mathbf{M}$ as accurately as possible.
@@
As a first motivating example, consider pointwise mutual information (PMI) constructed from different electronic healthcare records (EHR) datasets. PMI quantities the association between a pair of clinical concepts, and matrices representing these associations can be derived from co-occurrence summaries of various EHR datasets (Ahuja et al., 2020; Zhou et al., 2022). However, due to the lack of interoperability across healthcare systems (Rajkomar et al., 2018), different EHR data often involve non-identical concepts with limited overlap, resulting in substantial differences among their PMI matrices. The analysis of PMI matrices from different EHR datasets can thus be viewed as a multi-source matrix integration problem. Specifically, let $\mathcal{U}$ represent some concept set and suppose there is a symmetric PMI matrix $\mathbf{P}\in\mathbb{R}^{N\times N}$ associated with $\mathcal{U}$, where $N:=|\mathcal{U}|$. For the $i$th EHR, we denote its clinical concept by $\mathcal{U}_{i}\subset\mathcal{U}$ and let $n_{i}:=|\mathcal{U}_{i}|$. The PMI matrix derived from the $i$th EHR, $\mathbf{A}^{(i)}\in\mathbb{R}^{n_{i}\times n_{i}}$, then corresponds to the principal submatrix of $\mathbf{P}$ associated with $\mathcal{U}_{i}$. As it is often the case that the union of all the entries in $\{\mathbf{A}^{(i)}\}$ constitutes only a strict subset of those in $\mathbf{P}$, our aim is to integrate these $\{\mathbf{A}^{(i)}\}$ to recover the unobserved entries in $\mathbf{P}$.
@@
For another example, consider single-cell matrix data where rows represent genomic features, columns represent cells, and each entry records some specific information about a feature in the corresponding cell. A key challenge in the joint analysis for this type of data is to devise efficient computational strategies to integrate different data modalities (Ma et al., 2020; Lähnemann et al., 2020), as the experimental design may lead to a collection of single-cell data matrices for different, but potentially overlapping, sets of cells and features. More specifically, let $\mathbf{P}\in\mathbb{R}^{N\times M}$ be the population matrix for all involved features and cells where $N:=|\mathcal{U}|,M:=|\mathcal{V}|$ (with $\mathcal{U}$ and $\mathcal{V}$ denoting the sets of genomic features and cells, respectively). Each single-cell data matrix $\mathbf{A}^{(i)}\in\mathbb{R}^{n_{i}\times m_{i}}$ is then a submatrix of $\mathbf{P}$ corresponding to some $\mathcal{U}_{i}\subset\mathcal{U}$ and $\mathcal{V}_{i}\subset\mathcal{V}$; here we denote $n_{i}:=|\mathcal{U}_{i}|$ and $m_{i}:=|\mathcal{V}_{i}|$. Our aim is once again to integrate the collection of $\{\mathbf{A}^{(i)}\}$ to reconstruct the original $\mathbf{P}$.
@@
The above examples involving EHR and single-cell data are special cases of the matrix completion with noise and block-wise missing structures. However, the existing literature on matrix completion mainly focuses on recovering a possibly low-rank matrix based on uniformly sampled observed entries or independently sampled observed entries which may be contaminated by noise; see, e.g., Candès and Tao (2010); Candes and Recht (2012); Cai et al. (2010); Candes and Plan (2011); Koltchinskii et al. (2011); Tanner and Wei (2013); Chen et al. (2019); Fornasier et al. (2011); Mohan and Fazel (2012); Lee and Bresler (2010); Vandereycken (2013); Hu et al. (2012); Sun and Luo (2016); Cho et al. (2017); Chen et al. (2020); Yan et al. (2024); Srebro and Salakhutdinov (2010); Foygel et al. (2011); Cai and Zhou (2016); Keshavan et al. (2010) for an incomplete list of references.
@@
These assumptions of uniform or independent sampling in standard matrix completion models are generally violated in applications of matrix integration, thus necessitating the development of efficient methods for tackling the block-wise missing structures. Some examples of this development include the generalized integrative principal component analysis (GIPCA) of Zhu et al. (2020), structured matrix completion (SMC) of Cai et al. (2016), block-wise overlapping noisy matrix integration (BONMI) of Zhou et al. (2023), and symmetric positive semidefinite matrix completion (SPSMC) of Bishop and Yu (2014). The GIPCA procedure operates under the setting where each data matrix have some common samples and completely different variables, and furthermore assumes that each entry in these matrices are from some exponential family of distribution, with entries in the same matrix having the same distributional form. SMC is a spectral procedure for recovering the missing block of an approximately low-rank matrix when a subset of the rows and columns are observed; thus, SMC is designed to impute only a single missing block at a time. BONMI is also a spectral procedure for recovering a missing block (or submatrix) in an approximately low-rank matrix but, in contrast to SMC, assumes that this missing block is associated with a given pair of observed submatrices that share some (limited) overlap. SPSMC has a similar spectral procedure with BONMI to recover a low-rank symmetric positive semidefinite matrix using some observed principal submatrices. While BONMI combines submatrices pair by pair, SPSMC sequentially integrates each new submatrix with the combined structure formed by all previously integrated submatrices. The key idea behind BONMI and SPSMC is to align (via an orthogonal transformation) the spectral embeddings given by the leading (scaled) eigenvectors of the two overlapping submatrices and then impute the missing block by taking the outer product of these aligned embeddings. The use of embedding alignments also appeared in other applications including bilingual dictionary induction (Kementchedjhieva et al., 2018), knowledge graphs integration (Lin et al., 2019; Fanourakis et al., 2023), and vertex nominations (Zheng et al., 2022).
@@
In this paper, we extend the BONMI procedure, which handles only two overlapping submatrices, to $K\geq 2$ submatrices and propose the Chain-linked Multiple Matrix Integration (CMMI) for more efficient and flexible matrix completion. As a motivating example, suppose we have two overlapping pairs of submatrices $(\mathbf{A}^{(1)},\mathbf{A}^{(2)})$ and $(\mathbf{A}^{(2)},\mathbf{A}^{(3)})$. Using the overlapping entries between $\mathbf{A}^{(1)}$ and $\mathbf{A}^{(2)}$ (resp. $\mathbf{A}^{(2)}$ and $\mathbf{A}^{(3)}$) we can find an orthogonal transformation $\mathbf{W}^{(1,2)}$ (resp. $\mathbf{W}^{(2,3)}$) to align the embeddings $\hat{\mathbf{X}}^{(1)}$ and $\hat{\mathbf{X}}^{(2)}$ (resp. $\hat{\mathbf{X}}^{(2)}$ and $\hat{\mathbf{X}}^{(3)}$). Then by combining $\mathbf{W}^{(1,2)}$ and $\mathbf{W}^{(2,3)}$, we can also align $\hat{\mathbf{X}}^{(1)}$ to $\hat{\mathbf{X}}^{(3)}$ and recover the missing block associated with $\mathbf{A}^{(1)}$ and $\mathbf{A}^{(3)}$ even when these submatrices are non-overlapping. Generalizing this observation we can show that as long as $\{\mathbf{A}^{(i)}\}$ are connected then we can integrate them simultaneously to recover all the missing entries; here two submatrices $\mathbf{A}^{(i)}$ and $\mathbf{A}^{(j)}$ are said to be connected if there exists a sequence $i_{0},i_{1},\dots,i_{L}$ with $i_{0}=i$, $i_{L}=j$ such that $\mathbf{A}^{(i_{\ell-1})}$ and $\mathbf{A}^{(i_{\ell})}$ are overlapping for all $\ell=1,\dots,L$. The use of CMMI thus enables the recovery of many missing blocks that are unrecoverable by BONMI and furthermore allows for significantly smaller overlap between the observed submatrices. CMMI considers all possible overlapping pairs without relying on the integration order of submatrices, unlike SPSMC, enabling a more optimal recovery result.
@@
The structure of our paper is as follows. In Section 2 we introduce the model for multiple observed principal submatrices of a whole symmetric positive semi-definite matrix, and propose CMMI to integrate a chain of connected overlapping submatrices. Theoretical results for our CMMI procedures are presented in Section 3. In particular we derive error bounds in two-to-infinity norm for the spectral embedding of the submatrices and entrywise error bound for the recovered entries. Using these error bounds we show that our recovered entries are approximately normally distributed around their true values and that our algorithm yields consistent estimate even when there are only minimal overlap between the submatrices. We emphasize that the results in Section 3 also hold for BONMI (which is a special case of our results for $K=2$) and SPSMC, thereby providing significant refinements over those in Zhou et al. (2023) and Bishop and Yu (2014), which mainly focus on bounding the spectral or Frobenius norm errors of the missing block and embeddings. And our analysis handles both noisy and missing entries in the observed submatrices while Zhou et al. (2023) and Bishop and Yu (2014) only consider the case of noisy entries. Numerical simulations and experiments on real data are presented in Sections 4 and 5. In Section 6, we extend our embedding alignment approach to the cases of symmetric indefinite matrices and asymmetric or rectangular matrices. Detailed proofs of stated results are provided in the supplementary material. Section A in the supplementary material explores more complex matrix integration challenges, such as scenarios where the connected submatrices do not form a single chain but rather multiple chains with possibly quite different lengths and thus we need to select a suitably optimal chain among these candidates. The theoretical results in Section 3 allows us to develop several effective strategies for addressing these issues.
@@ 1.1 Notations
[…]
$\|\mathbf{M}\|_{2\to\infty}=\max_{\|\bm{x}\|=1}\|\mathbf{M}\bm{x}\|_{\infty}=\max_{i}\|\mathbf{m}_{i}\|,$
where $\mathbf{m}_{i}$ denotes the $i$th row of $\mathbf{M}$, i.e., $\|\mathbf{M}\|_{2\to\infty}$ is the maximum of the $\ell_{2}$ norms of the rows of $\mathbf{M}$. We note that the $2\to\infty$ norm is not sub-multiplicative. However, for any matrices $\mathbf{M}$ and $\mathbf{N}$ of conformal dimensions, we have
$\|\mathbf{M}\mathbf{N}\|_{2\to\infty}\leq\min\{\|\mathbf{M}\|_{2\to\infty}\times\|\mathbf{N}\|,\|\mathbf{M}\|_{\infty}\times\|\mathbf{N}\|_{2\to\infty}\};$
see Proposition 6.5 in Cape et al. (2019). Perturbation bounds using the $2\to\infty$ norm for the eigenvectors and/or singular vectors of a noisily observed matrix had recently attracted interests from the statistics community, see e.g., Chen et al. (2021); Cape et al. (2019); Fan et al. (2018); Abbe et al. (2020) and the references therein.

### v4

@@ 1 Introduction
The development of large-scale data collection and sharing has sparked considerable research interests in integrating data from diverse sources to efficiently uncover underlying signals. This problem is especially pertinent in fields such as healthcare research (Zhou et al., 2023), genomic data integration (Maneck et al., 2011; Tseng et al., 2015; Cai et al., 2016), single-cell data integration (Stuart et al., 2019; Ma et al., 2024), and chemometrics (Mishra et al., 2021). In this paper we consider a formulation of the problem where each source $i$ corresponds to a partially observed submatrix $\mathbf{M}^{(i)}$ of some matrix $\mathbf{M}$, and the goal is to integrate these $\{\mathbf{M}^{(i)}\}$ to recover $\mathbf{M}$ as accurately as possible.
@@
As a first motivating example, consider pointwise mutual information (PMI) constructed from different electronic healthcare records (EHR) datasets. PMI quantifies the association between a pair of clinical concepts, and matrices representing these associations can be derived from co-occurrence summaries of various EHR datasets (Ahuja et al., 2020; Zhou et al., 2022). However, due to the lack of interoperability across healthcare systems (Rajkomar et al., 2018), different EHR data often involve non-identical concepts with limited overlap, resulting in substantial differences among their PMI matrices. The analysis of PMI matrices from different EHR datasets can thus be viewed as a multi-source matrix integration problem. Specifically, let $\mathcal{U}$ represent some concept set and suppose there is a symmetric PMI matrix $\mathbf{P}\in\mathbb{R}^{N\times N}$ associated with $\mathcal{U}$, where $N:=|\mathcal{U}|$. For the $i$th EHR, we denote its clinical concept by $\mathcal{U}_{i}\subset\mathcal{U}$ and let $n_{i}:=|\mathcal{U}_{i}|$. The PMI matrix derived from the $i$th EHR, $\mathbf{A}^{(i)}\in\mathbb{R}^{n_{i}\times n_{i}}$, then corresponds to the principal submatrix of $\mathbf{P}$ associated with $\mathcal{U}_{i}$. As it is often the case that the union of all the entries in $\{\mathbf{A}^{(i)}\}$ constitutes only a strict subset of those in $\mathbf{P}$, our aim is to integrate these $\{\mathbf{A}^{(i)}\}$ to recover the unobserved entries in $\mathbf{P}$. Another example involving symmetric matrices integration appears in neuroscience, where symmetric covariance matrices are computed from calcium imaging data to characterize functional connectivity among neurons. Due to experimental constraints, only a strict subset of neurons are observed in each recording session. Integrating these incomplete covariance matrices enables the reconstruction of global neuronal interaction networks and accurate identification of functional hubs (Chang et al., 2022).
@@
An example of asymmetric matrix integration arises in single-cell matrix data, where rows represent genomic features, columns represent cells, and each entry records some specific information about a feature in the corresponding cell. A key challenge in the joint analysis for this type of data is to devise efficient computational strategies to integrate different data modalities (Ma et al., 2020; Lähnemann et al., 2020), as the experimental design may lead to a collection of single-cell data matrices for different, but potentially overlapping, sets of cells and features, such as those generated across batches, tissues, or technologies. Completing such partially overlapping data is crucial for constructing unified representations of cell populations and improving downstream tasks like clustering or trajectory inference. More specifically, let $\mathbf{P}\in\mathbb{R}^{N\times M}$ be the population matrix for all involved features and cells where $N:=|\mathcal{U}|,M:=|\mathcal{V}|$ (with $\mathcal{U}$ and $\mathcal{V}$ denoting the sets of genomic features and cells, respectively). Each single-cell data matrix $\mathbf{A}^{(i)}\in\mathbb{R}^{n_{i}\times m_{i}}$ is then a submatrix of $\mathbf{P}$ corresponding to some $\mathcal{U}_{i}\subset\mathcal{U}$ and $\mathcal{V}_{i}\subset\mathcal{V}$; here we denote $n_{i}:=|\mathcal{U}_{i}|$ and $m_{i}:=|\mathcal{V}_{i}|$. Our aim is once again to integrate the collection of $\{\mathbf{A}^{(i)}\}$ to reconstruct the original $\mathbf{P}$.
@@
The above examples involving EHR and single-cell data are special cases of the matrix completion with noise and block-wise missing structures. However, the existing literature on matrix completion mainly focuses on recovering a possibly low-rank matrix based on uniformly sampled observed entries or independently sampled observed entries which may be contaminated by noise; see, e.g., Candes and Recht (2012); Cai et al. (2010); Candes and Plan (2011); Koltchinskii et al. (2011); Tanner and Wei (2013); Chen et al. (2019); Fornasier et al. (2011); Mohan and Fazel (2012); Lee and Bresler (2010); Vandereycken (2013); Hu et al. (2012); Sun and Luo (2016); Cho et al. (2017); Chen et al. (2020); Srebro and Salakhutdinov (2010); Cai and Zhou (2016); Foygel et al. (2011) for an incomplete list of references.
@@
These assumptions of uniform or independent sampling in standard matrix completion models are generally violated in applications of matrix integration, thus necessitating the development of efficient methods for tackling the block-wise missing structures. Some examples of this development include the generalized integrative principal component analysis (GIPCA) of Zhu et al. (2020), structured matrix completion (SMC) of Cai et al. (2016), block-wise overlapping noisy matrix integration (BONMI) of Zhou et al. (2023), and symmetric positive semidefinite matrix completion (SPSMC) of Bishop and Yu (2014). The GIPCA procedure operates under the setting where each data matrix have some common samples and completely different variables, and furthermore assumes that each entry in these matrices are from some exponential family of distribution, with entries in the same matrix having the same distributional form. SMC is a spectral procedure for recovering the missing block of an approximately low-rank matrix when a subset of the rows and columns are observed; thus, SMC is designed to impute only a single missing block at a time. BONMI is also a spectral procedure for recovering a missing block (or submatrix) in an approximately low-rank matrix but, in contrast to SMC, assumes that this missing block is associated with a given pair of observed submatrices that share some (limited) overlap. SPSMC has a similar spectral procedure with BONMI to recover a low-rank symmetric positive semidefinite matrix using some observed principal submatrices. While BONMI combines submatrices pair by pair, SPSMC sequentially integrates each new submatrix with the combined structure formed by all previously integrated submatrices. The key idea behind BONMI and SPSMC is to align (via an orthogonal transformation) the spectral embeddings given by the leading (scaled) eigenvectors of the two overlapping submatrices and then impute the missing block by taking the outer product of these aligned embeddings.
@@
In this paper, we extend the BONMI procedure, which integrates only two overlapping submatrices at a time, to simultaneously and jointly integrate $K\geq 2$ submatrices, and propose the Chain-linked Multiple Matrix Integration (CMMI) for more efficient and flexible matrix completion. As a motivating example, suppose we have two overlapping pairs of positive semidefinite submatrices $(\mathbf{A}^{(1)},\mathbf{A}^{(2)})$ and $(\mathbf{A}^{(2)},\mathbf{A}^{(3)})$. For each submatrix $i$, let $\widehat{\mathbf{X}}^{(i)}$ be the $n_{i}\times d$ matrix whose columns are the $d$ leading eigenvectors of $\mathbf{A}^{(i)}$ scaled by the square root of the corresponding eigenvalues. The rows of $\widehat{\mathbf{X}}^{(i)}$ represent the embeddings of the $n_{i}$ entities associated with $\mathbf{A}^{(i)}$ into $\mathbb{R}^{d}$, and $\widehat{\mathbf{X}}^{(i)}\widehat{\mathbf{X}}^{(i)\top}$ correspond to the best rank-$d$ approximation to $\mathbf{A}^{(i)}$. However, as the leading eigenvectors of $\mathbf{A}^{(i)}$ are not necessarily unique, we cannot directly use the inner product between different $\widehat{\mathbf{X}}^{(i)}$ to estimate the unobserved entries. Rather, we first have to align $\{\widehat{\mathbf{X}}^{(1)},\widehat{\mathbf{X}}^{(2)},\widehat{\mathbf{X}}^{(3)}\}$ using their overlapping submatrices. More specifically, we align $\widehat{\mathbf{X}}^{(1)}$ and $\widehat{\mathbf{X}}^{(2)}$ by finding the orthogonal transformation $\mathbf{W}^{(1,2)}$ that maps the embeddings of $\mathcal{U}_{1}\cap\mathcal{U}_{2}$ in $\widehat{\mathbf{X}}^{(1)}$ to (approximate) their counterparts in $\widehat{\mathbf{X}}^{(2)}$. The entries of $\widehat{\mathbf{X}}^{(1)}\mathbf{W}^{(1,2)}\widehat{\mathbf{X}}^{(2)\top}$ then serve as estimates of the unobserved entries between $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$. Similarly, we compute $\mathbf{W}^{(2,3)}$ to map the embeddings for $\mathcal{U}_{2}\cap\mathcal{U}_{3}$ in $\widehat{\mathbf{X}}^{(2)}$ to their counterparts in $\widehat{\mathbf{X}}^{(3)}$, and entries of $\widehat{\mathbf{X}}^{(2)}\mathbf{W}^{(2,3)}\widehat{\mathbf{X}}^{(3)\top}$ serve as estimates of the unobserved entries between $\mathcal{U}_{2}$ and $\mathcal{U}_{3}$. Finally, we can use $\widehat{\mathbf{X}}^{(1)}\mathbf{W}^{(1,2)}\mathbf{W}^{(2,3)}\widehat{\mathbf{X}}^{(3)\top}$ to estimate the unobserved entries between $\mathcal{U}_{1}$ and $\mathcal{U}_{3}$ even when $\mathcal{U}_{1}\cap\mathcal{U}_{3}=\varnothing$ (so that $\mathbf{A}^{(1)}$ and $\mathbf{A}^{(3)}$ are non-overlapping). Generalizing this observation we can show that as long as $\{\mathbf{A}^{(i)}\}$ are connected then we can integrate them simultaneously to recover all the missing entries; here two submatrices $\mathbf{A}^{(i)}$ and $\mathbf{A}^{(j)}$ are said to be connected if there exists a sequence $i_{0},i_{1},\dots,i_{L}$ with $i_{0}=i$, $i_{L}=j$ such that $\mathbf{A}^{(i_{\ell-1})}$ and $\mathbf{A}^{(i_{\ell})}$ are overlapping for all $\ell=1,\dots,L$. The use of CMMI thus enables the recovery of many missing blocks that are unrecoverable by BONMI and furthermore allows for significantly smaller overlap between the observed submatrices. CMMI considers all possible overlapping pairs without relying on the integration order of submatrices, unlike SPSMC, enabling a more optimal recovery result.
@@
The structure of our paper is as follows. In Section 2 we introduce the model for multiple observed principal submatrices of a whole symmetric positive semi-definite matrix, and propose CMMI to integrate a chain of connected overlapping submatrices. Theoretical results for our CMMI procedures are presented in Section 3. In particular we derive error bounds in $2\to\infty$ norm for the spectral embedding of the submatrices and entrywise error bound for the recovered entries. Using these error bounds we show that our recovered entries are approximately normally distributed around their true values and that CMMI yields consistent estimates even with minimal overlaps between the observed submatrices. We emphasize that the results in Section 3 also hold for BONMI (which is a special case of our results for $K=2$) and SPSMC, thereby providing significant refinements over those in Zhou et al. (2023) and Bishop and Yu (2014), which mainly focus on bounding the spectral or Frobenius norm errors of the missing block and embeddings. And our analysis handles both noisy and missing entries in the observed submatrices while Zhou et al. (2023) and Bishop and Yu (2014) only consider the case of noisy entries. Numerical simulations and experiments on real data are presented in Sections 4 and 5. In Section 6, we extend our embedding alignment approach to the cases of symmetric indefinite matrices and asymmetric or rectangular matrices. Detailed proofs of stated results and additional numerical experiments are provided in the supplementary material. Section B of the supplementary material extends the basic CMMI algorithm for chains to handle the integration of connected submatrices with arbitrarily complex structures, making it more applicable to real-world scenarios. We provide theoretical analysis of the generalized CMMI based on the results in Section 3, and demonstrate its effectiveness through real data experiments.
@@ 1.1 Notations
We summarize some notations used in this paper. For any positive integer $n$, we denote by $[n]$ the set $\{1,2,\dots,n\}$. For two non-negative sequences $\{a_{n}\}_{n\geq 1}$ and $\{b_{n}\}_{n\geq 1}$, we write $a_{n}\lesssim b_{n}$ (resp. $a_{n}\gtrsim b_{n}$) if there exists some constant $C>0$ such that $a_{n}\leq Cb_{n}$ (resp. $a_{n}\geq Cb_{n}$) for all $n\geq 1$, and we write $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $a_{n}\gtrsim b_{n}$. The notation $a_{n}\ll b_{n}$ (resp. $a_{n}\gg b_{n}$) means that there exists some sufficiently small (resp. large) constant $C>0$ such that $a_{n}\leq Cb_{n}$ (resp. $a_{n}\geq Cb_{n}$). If $a_{n}/b_{n}$ stays bounded away from $\infty$, we write $a_{n}=O(b_{n})$ and $b_{n}=\Omega(a_{n})$, and we use the notation $a_{n}=\Theta(b_{n})$ to indicate that $a_{n}=O(b_{n})$ and $a_{n}=\Omega(b_{n})$. If $a_{n}/b_{n}\to 0$, we write $a_{n}=o(b_{n})$ and $b_{n}=\omega(a_{n})$. We say a sequence of events $\mathcal{A}_{n}$ holds with high probability if for any $c>0$ there exists a finite constant $n_{0}$ depending only on $c$ such that $\mathbb{P}(\mathcal{A}_{n})\geq 1-n^{-c}$ for all $n\geq n_{0}$. We write $a_{n}=O_{p}(b_{n})$ (resp. $a_{n}=o_{p}(b_{n})$) to denote that $a_{n}=O(b_{n})$ (resp. $a_{n}=o(b_{n})$) holds with high probability. We denote by $\mathcal{O}_{d}$ the set of $d\times d$ orthogonal matrices. For any matrix $\mathbf{M}\in\mathbb{R}^{A\times B}$ and index sets $\mathcal{A}\subseteq[A]$, $\mathcal{B}\subseteq[B]$, we denote by $\mathbf{M}_{\mathcal{A},\mathcal{B}}\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{B}|}$ the submatrix of $\mathbf{M}$ formed from rows $\mathcal{A}$ and columns $\mathcal{B}$, and we denote by $\mathbf{M}_{\mathcal{A}}\in\mathbb{R}^{|\mathcal{A}|\times B}$ the submatrix of $\mathbf{M}$ consisting of the rows indexed by $\mathcal{A}$. The Hadamard product between conformal matrices $\mathbf{M}$ and $\mathbf{N}$ is denoted by $\mathbf{M}\circ\mathbf{N}$. Given a matrix $\mathbf{M}$, we denote its spectral, Frobenius, and infinity norms by $\|\mathbf{M}\|$, $\|\mathbf{M}\|_{F}$, and $\|\mathbf{M}\|_{\infty}$. We also denote the maximum entry (in modulus) of $\mathbf{M}$ by $\|\mathbf{M}\|_{\max}$ and the $2\to\infty$ norm of $\mathbf{M}$ by $\|\mathbf{M}\|_{2\to\infty}=\max_{\|\bm{x}\|=1}\|\mathbf{M}\bm{x}\|_{\infty}=\max_{i}\|\mathbf{m}_{i}\|,$ where $\mathbf{m}_{i}$ denotes the $i$th row of $\mathbf{M}$, i.e., $\|\mathbf{M}\|_{2\to\infty}$ is the maximum of the $\ell_{2}$ norms of the rows of $\mathbf{M}$.